### v1

# Locally-APN Binomials with Low Boomerang Uniformity in Odd Characteristic

Abstract
Recently, several studies have shown that when $q\equiv 3\pmod{4}$, the function $F_{r}(x)=x^{r}+x^{r+\frac{q-1}{2}}$ defined over $\mathbb{F}_{q}$ is locally-APN and has boomerang uniformity at most $2$. In this paper, we extend these results by showing that if there is at most one $x\in\mathbb{F}_{q}$ with $\chi(x)=\chi(x+1)=1$ satisfying $(x+1)^{r}-x^{r}=b$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $F_{r}$ is locally-APN with boomerang uniformity at most $2$. Moreover, we study the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$, and the boomerang spectrum of $F_{2}$ when $p=3$.
Keywords. Binomials, Differential Uniformity, Differential Spectrum, Boomerang Uniformity, Boomerang Spectrum, Locally APN Functions.
Mathematics Subject Classification(2020) 94A60, 06E30
1 Introduction
In this paper, let $q=p^{n}$ be an odd prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. In recent years, considerable attention has been devoted to the study of the differential properties of functions of the form
$F_{r,u}(x)=x^{r}(1+u\chi(x)),$ (1)
where $\chi(x)=x^{\frac{q-1}{2}}$ is the quadratic character over $\mathbb{F}_{q}$. Differential uniformity, introduced by Nyberg[16], quantifies the resistance of an S-box in a block cipher to differential cryptanalysis. The definition of the differential uniformity is given below.
Definition 1.1.
Let $F$ be a function on $\mathbb{F}_{q}$. Define $\delta_{F}(a,b)$ by the number of solutions of F(x+a)F(x)=bF(x+a)-F(x)=b, where $a\in\mathbb{F}_{q}^{*}$ and $b\in\mathbb{F}_{q}$. Then, the differential uniformity of $F$ is defined by :
$\delta_{F}=\displaystyle\max_{a\in\mathbb{F}_{q}^{*},b\in\mathbb{F}_{q}}\delta_{F}(a,b).$
A function on $\mathbb{F}_{q}$ is said to be perfect nonlinear (PN), if its differential uniformity is equal to $1$. When its differential uniformity equals $2$, the function is called almost perfect nonlinear (APN). In [15], Ness and Helleseth showed that $F_{r,u}$ is APN when r=q2r=q-2 and $\chi(u-1)=\chi(u+1)=\chi(u)$ for $p=3$. Subsequently, Zeng et al. [21] extended this result to the case $q\equiv 3\pmod{4}$, proving that $F_{q-2,u}$ is APN if $\chi(u-1)=\chi(u+1)=-\chi(5u\pm 3)$.
More recently, several studies have focused on the non-APN case of $F_{q-2,u}$. For instance, Xia et al. [18] investigated the differential uniformity of $F_{q-2,u}$ and determined its differential spectrum when $\chi(u+1)=\chi(u-1)$ in characteristic $3$. Their results were later extended in [19] to the general case where $q\equiv 3\pmod{4}$. Ren et al. [17] studied the differential spectrum of $F_{q-2,u}$ in the complementary case $\chi(u+1)\neq\chi(u-1)$, while Lyn et al. [13] independently showed that $F_{q-2,u}$ is a differentially $4$-uniform permutation under the same condition.
Budaghyan and Pal [3] conjectured, based on experimental evidence, that there might exist an infinite subclass of APN functions among $F_{2,u}$. However, this conjecture was disproved in [14], where the authors conducted a detailed study of the differential uniformity of $F_{2,u}$. The same conjecture was independently disproved in [1] by means of function field theory. Koo and Kwon [10] further investigated the differential uniformity of $F_{r,u}$ for $r=\frac{q+1}{4}$. The differential properties of $F_{r,u}$ are summarized in Table 1.
$r$ $u$ Conditions $\delta_{F_{r,u}}$ Spectrum Ref.
q2q-2 $\pm 1$ $p=3$, $n>3$ : odd $\frac{q+1}{4}$ (locally-PN) O [13]
q2q-2 $u\neq\pm 1$ $p=3$, $n>3$ : odd $\chi(1+u)=\chi(u-1)=\chi(u)$ 2 (APN) O [18]
q2q-2 $u\neq\pm 1$ $p=3$, $n>3$ : odd $\chi(1+u)=\chi(u-1)\neq\chi(u)$ 33 O [18]
q2q-2 $u\neq\pm 1$ $p=3$, $n>3$ : odd $\chi(1+u)\neq\chi(u-1)$ $\delta_{F_{r,u}}\leq 4$ O [17]
q2q-2 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $\delta_{F_{r,u}}\leq 5$ X [3]
22 $\pm 1$ $q\equiv 3\pmod{4}$ $\frac{q+1}{4}$ (locally-APN) O [14]
22 $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $3\leq\delta_{F_{r,u}}\leq 5$ O [14]
$\frac{q+1}{4}$ $u\neq\pm 1$ $q\equiv 3\pmod{4}$ $\delta_{F_{r,u}}\leq 5$ X [10]
$\frac{q+1}{4}$ $\pm 1$ $q\equiv 3\pmod{8}$ $\frac{q+1}{4}$ (locally-PN) O [10]
$\frac{q+1}{4}$ $\pm 1$ $q\equiv 7\pmod{8}$ $\frac{q+1}{4}$ (locally-APN) O [10]
$p^{k}+1$ $\pm 1$ $n$ : odd $\frac{q+1}{4}$ (locally-APN) X This paper
$\frac{3^{k}+1}{2}$ $\pm 1$ $p=3$, $n$ : odd $\frac{q+1}{4}$ (locally-APN) X This paper
33 $\pm 1$ $q\equiv 11\pmod{12}$ $\frac{q+1}{4}$ (locally-APN) O This paper
$\frac{2q-1}{3}$ $\pm 1$ $q\equiv 11\pmod{12}$ $\frac{q+1}{4}$ (locally-APN) O This paper
$\frac{3^{\frac{n+1}{2}}-1}{2}$ $\pm 1$ p=3,np=3,\;n : odd $\frac{q+1}{4}$ (locally-APN) X This paper
$\frac{3^{n+1}-1}{8}$ $\pm 1$ p=3,np=3,\;n : odd $\frac{q+1}{4}$ (locally-APN) X This paper
Table 1: Differential uniformity of $F_{r,u}$
The boomerang uniformity was introduced as the maximum value of the boomerang connectivity table (BCT), a tool proposed to analyze the boomerang attack, which is a variant of differential cryptanalysis [4]. Originally, this notion was defined only for permutations over finite fields of characteristic $2$. Later, Li et al. [11] extended it to general functions, not necessarily permutations, so that it can be applied to mappings in arbitrary characteristic.
Definition 1.2.
Let $F$ be a function on $\mathbb{F}_{q}$. Define $\beta_{F}(a,b)$ by the number of common solutions $(x,y)$ of the following system :
$\begin{cases}F(x)-F(y)=b,\\ F(x+a)-F(y+a)=b.\end{cases}$
Then the boomerang uniformity of $F$ is given by :
$\beta_{F}=\displaystyle\max_{a,b\in\mathbb{F}_{q}^{*}}\beta_{F}(a,b).$
For the case $u\neq\pm 1$, the existing studies have focused only on the differential properties of $F_{r,u}$, and no results on their boomerang behavior have been reported so far. In contrast, when $u=\pm 1$, both the differential and boomerang properties of $F_{r,u}$ have been investigated in several recent works. In the following, we concentrate on this particular case. Note that $F_{r,-1}$ is affine equivalent to $F_{r,1}$. In this case, we simply write
$F_{r}(x)=x^{r}(1+\chi(x))=x^{r}+x^{r+\frac{q-1}{2}}.$
Recently, several studies have investigated the differential and boomerang spectra of $F_{r}$ for various choices of $r$. These works consistently show that $F_{r}$ is locally-APN and has boomerang uniformity at most $2$ in all known cases. Lyn et al. [13] determined the differential and boomerang spectra of $F_{q-2}$, showing that it is locally-PN with boomerang uniformity 0 when $p=3$, and locally-APN with boomerang uniformity at most $1$ when $p\neq 3$. Mesnager and Wu [14] proved that $F_{2}$ is locally-APN with boomerang uniformity at most $2$, and further investigated its differential spectrum. They also proved that the boomerang uniformity of $F_{2}$ equals $2$, when $q\geq 9613^{2}$. However, we will show in Section 4.2 that the boomerang uniformity of $F_{2}$ equals $1$ when $p=3$. Yan and Ren [20] independently analyzed the differential spectrum of $F_{2}$ using a different approach. Koo and Kwon [10] studied $F_{r}$ for $r=\frac{q+1}{4}$, showing that $F_{r}$ is locally-PN with boomerang uniformity 0 when $q\equiv 3\pmod{8}$, and locally-APN with boomerang uniformity $2$ when $q\equiv 7\pmod{8}$. They also provided complete characterizations of the differential and boomerang spectra of $F_{r}$ in these cases. The boomerang properties of $F_{r}$ are summarized in Table 2.
$r$ Conditions $\beta_{F_{r}}$ Spectrum Ref.
q2q-2 $p=3$ 0 - [13]
q2q-2 $q\equiv 3\pmod{4}$, $p>3$ 11 O [13]
22 $q\equiv 3\pmod{4}$, large $q$ 2 X [14]
$\frac{q+1}{4}$ $q\equiv 3\pmod{8}$ 0 - [10]
$\frac{q+1}{4}$ $q\equiv 7\pmod{8}$ 2 O [10]
$p^{k}+1$ $n$ : odd $\leq 2$ X This paper
$\frac{3^{k}+1}{2}$ p=3,np=3,\;n : odd $\leq 2$ X This paper
33 $q=11\;(\bmod 12)$ $\leq 2$ X This paper
$\frac{2q-1}{3}$ $q=11\;(\bmod 12)$ $\leq 2$ X This paper
$\frac{3^{n+1}-1}{2}$ p=3,np=3,\;n : odd $\leq 2$ X This paper
$\frac{3^{n+1}-1}{8}$ p=3,np=3,\;n : odd $\leq 2$ X This paper
22 $p=3,\;n\geq 3$ : odd 1 O This paper
Table 2: Boomerang uniformity of $F_{r}$.
In this paper, we study the differential and boomerang properties of $F_{r}$ when $q\equiv 3\pmod{4}$. In particular, we show that if the equation $(x+1)^{r}-x^{r}=b$ has at most one solution satisfying $\chi(x)=\chi(x+1)=1$ and $\gcd(r,q-1)\mid 2$, then $\delta_{F_{r}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$. This implies that $F_{r}$ is locally-APN. Furthermore, we also prove that its boomerang uniformity is at most $2$. The above condition is more likely to hold when the differential uniformity of $x^{r}$ is small, and it certainly holds when $x^{r}$ is PN. In this paper, we show that this is indeed the case for the exponents $r$ listed in Table 3 in Section 2. In addition, we further investigate the differential spectrum for $r=3$ and $r=\frac{2q-1}{3}$, and the boomerang spectrum of $F_{2}$ when $p=3$. By Lemma 2.8, the latter result can be extended to the case k=n1k=n-1 for $r=\frac{3^{k}+1}{2}$.
The rest of this paper is organized as follows. Section 2 provides the necessary preliminaries for our results. In Section 3, we investigate the differential properties of $F_{r}$. More precisely, for our exponents $r$ in Table 3 we prove that $F_{r}$ is locally-APN in Section 3.1, and study the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$ in Sections 3.2 and 3.3, respectively. Section 4 focuses on the boomerang properties of $F_{r}$: we show that for our exponents $r$ in Table 3 $F_{r}$ has boomerang uniformity at most $2$ in Section 4.1, and determine the boomerang spectrum of $F_{2}$ for $p=3$ in Section 4.2. Finally, Section 5 concludes the paper.
2 Preliminaries
When $F(x)=x^{r}$ is a power function, multiplying $\frac{1}{a^{r}}$ on the both sides of $b=F(x+a)-F(x)=(x+a)^{d}-x^{d}$, we have
$\frac{b}{a^{r}}=\left(\displaystyle\frac{x}{a}+1\right)^{r}-\left(\displaystyle\frac{x}{a}\right)^{r}=(y+1)^{r}-y^{r},$
where $y=\displaystyle\frac{x}{a}$. Thus, we have
$\delta_{F}(a,b)=\delta_{F}\left(1,\displaystyle\frac{b}{a^{r}}\right),\text{ so }\delta_{F}=\displaystyle\max_{b\in\mathbb{F}_{q}}\delta_{F}(1,b),$ (2)
in this case. Notion of locally-APN functions was introduced in [2], for the even characteristic case. In [9], Hu et al generalized it for odd characteristic cases, as follows.
$\delta_{F}(1,b)\leq 2\text{ for all }b\in\mathbb{F}_{q}\setminus\mathbb{F}_{p}.$ (3)
In our previous paper [10], we extend the above definition to functions with the property
$\delta_{F}(a,b)=\delta_{F}(1,g_{a}(b))\text{ for all }b\in\mathbb{F}_{p^{n}},$ (4)
where $g_{a}$ permutes $\mathbb{F}_{q}$ for each $a\in\mathbb{F}_{q}^{*}$.
Definition 2.1.
Let $F$ be a function on $\mathbb{F}_{q}$ satisfying (4) for every $a\in\mathbb{F}_{p^{n}}^{*}$. Then, $F$ is called locally-APN if $\delta_{F}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}\setminus\mathbb{F}_{p}$.
Lemma 2.2.
(Lemma 11 of [14]) Let $a\in\mathbb{F}_{q}^{*}$ and $b\in\mathbb{F}_{q}$. Then,
$\delta_{F_{r,u}}(a,b)$ $=\begin{cases}\delta_{F_{r,u}}\left(1,\frac{b}{a^{r}}\right)&\text{ if }\chi(a)=1,\\ \delta_{F_{r,u}}\left(1,\frac{b}{(-1)^{r+1}a^{r}}\right)&\text{ if }\chi(a)=-1,\end{cases}$
$\beta_{F_{r,u}}(a,b)$ $=\begin{cases}\beta_{F_{r,u}}\left(1,\frac{b}{a^{r}}\right)&\text{ if }\chi(a)=1,\\ \beta_{F_{r,u}}\left(1,\frac{b}{(-1)^{r}a^{r}}\right)&\text{ if }\chi(a)=-1.\end{cases}$
We denote
$S_{ij}=\{x\in\mathbb{F}_{q}:\chi(x)=(-1)^{i},\chi(x+1)=(-1)^{j}\},$
where $i,j\in\{0,1\}$. The following lemma is well-known and will be used in Section 3.2.
Lemma 2.3.
[6] If $p^{n}\equiv 1\pmod{4}$, then $S_{00}=\frac{p^{n}-5}{4}$ and $S_{01}=\#S_{10}=\#S_{11}=\frac{p^{n}-1}{4}$. If $p^{n}\equiv 3\pmod{4}$, then $S_{00}=\#S_{10}=\#S_{11}=\frac{p^{n}-3}{4}$ and $S_{01}=\frac{p^{n}+1}{4}$.
The following four lemmas on the quadratic character are useful for deriving the differential spectrum of $F_{3}$ in Section 3.2 and the boomerang spectrum of $F_{2}$ (for $p=3$) in Section 4.2.
Lemma 2.4.
(Theorem 5.48 of [12]) Let $f(x)=a_{2}x^{2}+a_{1}x+a_{0}\in\mathbb{F}_{p^{n}}[x]$ with $p$ odd and $a_{2}\neq 0$. Put $d=a_{1}^{2}-4a_{0}a_{2}$. Then,
$\sum_{x\in\mathbb{F}_{p^{n}}}\chi(f(x))=\begin{cases}-\chi(a_{2})&\text{ if }d\neq 0,\\ (p^{n}-1)\chi(a_{2})&\text{ if }d=0.\end{cases}$
Lemma 2.5.
Let $q\equiv 3\pmod{4}$ and $f$ be a function on $\mathbb{F}_{q}$ with f(x)=f(x)f(-x)=-f(x) for all $x\in\mathbb{F}_{q}$. Then, $\sum_{x\in\mathbb{F}_{q}}\chi(f(x))=0$.
Proof.
Applying y=xy=-x, we have
$\sum_{x\in\mathbb{F}_{q}}\chi(f(x))=\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi(f(-y))=\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi(-f(y))=-\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi(f(y)),$
which completes the proof. ∎
By the above lemma, we get $\sum_{x\in\mathbb{F}_{q}}\chi\left(x(x^{2}-1)\right)=0$, which will be used in Section 4.2.
Lemma 2.6.
(Theorem 5.41 of [12]) Let $\eta(\cdot)$ be a multiplicative character of $\mathbb{F}_{q}$ of order $m>1$ and let $f\in\mathbb{F}_{q}[x]$ be a monic polynomial of positive degree that is not $m$-th power of a polynomial. Let $d$ be the number of distinct roots of $f$ in its splitting field over $\mathbb{F}_{q}$. Then for every $a\in\mathbb{F}_{p^{n}}$ we have
$\left|\sum_{x\in\mathbb{F}_{q}}\eta(af(x))\right|\leq(d-1)\sqrt{q}.$
The following lemma is well-known and useful for our result.
Lemma 2.7.
(Lemma 9 of [8]) Let $p$, $k$, $n$ be positive integers with $k\leq n$. If $\frac{n}{\gcd(k,n)}$ is odd, then $\gcd(p^{k}+1,p^{n}-1)=2$.
In this paper, we show that for $r$ in Table 3 $F_{r}$ is locally-APN with differential uniformity $\frac{q+1}{4}$ with boomerang uniformity at most $2$. These exponents $r$ form a subclass of those satisfying $\gcd(r,q-1)=2$ for which the monomial $x^{r}$ has low differential uniformity. Note that the differential properties summarized in Table 3 are stated for the monomial $x^{r}$, but in some cases the equivalent form $x^{r+\frac{q-1}{2}}=x^{r}\chi(x)$ is actually PN or APN. Since $F_{r+\frac{q-1}{2}}(x)=x^{r+\frac{q-1}{2}}(1+\chi(x))=x^{r}(\chi(x)+1)=F_{r}(x)$ for all $x\in\mathbb{F}_{q}$, the two forms are identical as functions.
$r$ Conditions DU of $x^{r}$ Ref.
$p^{k}+1$ $n$ : odd PN [5]
$\frac{3^{k}+1}{2}$ $p=3$, $n$ : odd PN [5]
33 $q\equiv 11\pmod{12}$ APN -
$\frac{2q-1}{3}$ $q\equiv 11\pmod{12}$ APN Inverse of $x^{3}$
$\frac{3^{\frac{n+1}{2}}-1}{2}$ $p=3$, $n$ : odd APN [7]
$\frac{3^{n+1}-1}{8}$ $p=3$, $n$ : odd APN [7]
Table 3: List of $r$ and Differential Properties of $x^{r}$.
Moreover, when $r=\frac{3^{k}+1}{2}$ (the Coulter–Matthews exponent), $x^{r}$ is PN only for odd $k$ [5]. However, according to the linear equivalence established in Lemma 2.8, our results hold for all $k$. Two functions $F$ and $F^{\prime}$ defined on $\mathbb{F}_{q}$ are said to be linearly(affine) equivalent if there exist linear(affine) permutations $L_{1}$ and $L_{2}$ over $\mathbb{F}_{q}$ such that $F^{\prime}=L_{1}\circ F\circ L_{2}$. It is well known that the differential and boomerang spectra are invariant under affine equivalence.
Lemma 2.8.
Let $p=3$ and $n$ be odd. Then $F_{\frac{3^{k}+1}{2}}$ is linearly equivalent to $F_{\frac{3^{n-k}+1}{2}}$.
Proof.
Let $L(x)=x^{3^{n-k}}$. Then, for nonzero $x\in\mathbb{F}_{q}$,
$(F_{\frac{3^{k}+1}{2}}\circ L)(x)$ $=\left(x^{3^{n-k}}\right)^{\frac{3^{k}+1}{2}}(1+\chi(x^{3^{n-k}}))=x^{\frac{3^{n}+3^{n-k}}{2}}(1+\chi(x))=x^{\frac{3^{n-k}+1}{2}}\cdot x^{\frac{3^{n}-1}{2}}(1+\chi(x))$
$=x^{\frac{3^{n-k}+1}{2}}(\chi(x)+1)=F_{\frac{3^{n-k}+1}{2}}(x).$
Clearly, we can easily check that $(F_{\frac{3^{k}+1}{2}}\circ L)(0)=F_{\frac{3^{n-k}+1}{2}}(0)=0$. ∎
3 Differential Properties
In this section, we study the differential properties of $F_{r}$. We desire to count the number of solutions of
$b=F_{r}(x+1)-F_{r}(x)=(x+1)^{r}(1+\chi(x+1))-x^{r}(1+\chi(x)).$ (5)
Denote $D_{ij}(b)$ be the number of solutions of (5) in $S_{ij}$ where $i,j\in\{0,1\}$. Then the following is almost clear.
$\delta_{F_{r}}(1,b)=\begin{cases}\sum D_{ij}(b)&\text{ if }b\neq 0,2,\\ \sum D_{ij}(b)+1&\text{ if }b=0\text{ or }2.\end{cases}$
3.1 Differential Uniformity of $F_{r}$
In this subsection, we show that $F_{r}$ is locally-APN for the exponents $r$ listed in Table 3. First, we state the main theorem.
Theorem 3.1.
Let $r>1$ be an integer and $q=p^{n}\equiv 3\pmod{4}$ If $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $\delta_{F_{r}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$ and
$\delta_{F_{r}}=\delta_{F_{r}}(1,0)=\frac{q+1}{4},$ (6)
and hence $F_{r}$ is locally-APN. In particular, for each parameter set $(p,n,r)$ satisfying the corresponding conditions in Table 3, $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$, and therefore, $F_{r}$ is locally-APN with (6).
Note that if $x^{r}$ is PN, then the assumption of the above theorem holds clearly. The case $r=3$ will be discussed separately in Section 3.2, where we investigate the differential spectrum of $F_{3}$. We also find the differential spectrum of the case $r=\frac{2q-1}{3}$ in Section 3.3. To prove this theorem for the remaining cases, we investigate each $D_{ij}(b)$.
Lemma 3.2.
If $q\equiv 3\pmod{4}$, then
$D_{11}(b)=\begin{cases}\frac{p^{n}-3}{4},&\text{ if }b=0,\\ 0,&\text{ otherwise.}\end{cases}$
Proof.
If $x\in S_{11}$, then (5) implies $b=0$. By Lemma 2.3, we obtain $S_{11}=\frac{p^{n}-3}{4}$, as desired. ∎
Lemma 3.3.
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}$. If $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$, then $D_{00}(b)\leq 1$. If $\gcd(r,q-1)\in\{1,2\}$, then $D_{00}(0)=0$.
Proof.
If there is a solution $x\in S_{00}$, then we have
$(x+1)^{r}-x^{r}=\frac{b}{2}$ (7)
By assumption, we can easily see that $D_{00}(b)\leq 1$. If $x^{r}$ is PN, then there is at most one such $x$ satisfying (7), we have $D_{00}(b)\leq 1$.
If there is a solution $x\in S_{00}$ in (5) with $b=0$, then we have $(x+1)^{r}=x^{r}$. If $\gcd(r,q-1)=1$, then we have $x=x+1$, a contradiction. If $\gcd(r,q-1)=2$, then we obtain $x^{2}=(x+1)^{2}$, which implies $x=-\frac{1}{2}$. Then, we get $\chi(x)=\chi\left(-\frac{1}{2}\right)=\chi\left(\frac{1}{2}\right)=\chi(x+1)=1$, a contradiction. Therefore, $D_{00}(0)=0$. ∎
Lemma 3.4.
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}$. Then, $D_{01}(b)=0$ or $D_{10}(b)=0$.
Proof.
If there is a solution $x\in S_{01}$ of (5), then $x^{r}=-\frac{b}{2}$. Since $x\in S_{01}$, we have $\chi\left(\frac{b}{2}\right)=-1$. By the way, if there is a solution $x^{\prime}\in S_{10}$ of (5), then $(x^{\prime}+1)^{r}=\frac{b}{2}$. Since $x^{\prime}\in S_{10}$, we obtain $\chi\left(\frac{b}{2}\right)=-1$. Since $\chi\!\left(\frac{b}{2}\right)=1$ and $\chi\!\left(\frac{b}{2}\right)=-1$ cannot occur simultaneously, we conclude that $D_{01}(b)=0$ or $D_{10}(b)=0$. ∎
Lemma 3.5.
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}$. If $\gcd(r,q-1)\in\{1,2\}$, then $D_{01}(b)\leq 1$ and $D_{10}(b)\leq 1$. Furthermore, if $b\in\{0,2\}$, $D_{01}(b)=D_{10}(b)=0$.
Proof.
If there is a solution $x\in S_{01}$ of (5), then (5) reduces to $x^{r}=-\frac{b}{2}$. If $\gcd(r,q-1)=1$, then there is at most one $x\in\mathbb{F}_{q}$ satisfying $x^{r}=-\frac{b}{2}$. If $\gcd(r,q-1)=2$, then there are at most $2$ solutions of $x^{r}=-\frac{b}{2}$. In this case, $r$ is even, and hence if $x=\theta$ is a solution of $x^{r}=-\frac{b}{2}$, then the other solution is $x=-\theta$. Since $\chi(\theta)\neq\chi(-\theta)$, at most one of $\theta$ and $\theta$ belongs to $S_{01}$.
When $b=0$, $x^{r}=-\frac{b}{2}=0$ implies $x=0$, and hence there is no solution in $S_{01}$. When $b=2$, if there is $x\in S_{01}$ satisfying $x^{r}=-\frac{b}{2}=-1$, then $\chi(x^{r})=\left(\chi(x)\right)^{r}=1=\chi(-1)$, a contradiction. Therefore, $D_{01}(0)=D_{01}(2)=0$.
By symmetry, the other proofs are similar with above, and we omit here. ∎
We have now analyzed all the cases of $D_{ij}(b)$, and are ready to prove the main theorem of this subsection for all $r$ except $r=3$ and $r=\frac{2q-1}{3}$. It is worth noting that the lemmas established here also hold for these two exponents, and hence we have (8), which will be proved in Lemma 3.6. These lemmas will also be used in the following subsections, where we investigate the differential spectra of $F_{r}$ for $r=3$ and $r=\frac{2q-1}{3}$.
Proof of Theorem 3.1.
If we substitute $x=0$ and x=1x=-1 to (5), then we have $b=2$ and $b=0$, respectively. Thus, by lemmas in this subsection, we can see that $\delta_{F_{r}}(1,0)=\displaystyle\frac{q+1}{4}$, and
$\delta_{F_{r}}(1,b)\leq D_{00}(b)+1,\ \ \text{ if }b\in\mathbb{F}_{q}^{*}.$ (8)
It is clear that $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$, when $x^{r}$ is PN. It remains to show that $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$ when $x^{r}$ is APN, that is, for $r\in\left\{\frac{3^{\frac{n+1}{2}}-1}{2},\ \frac{3^{n+1}-1}{8}\right\}$, where $p=3$ and $n$ is odd. In [7], where the APN property of $x^{r}$ was proved, the key idea was to show that the equation $(x+1)^{r}-x^{r}=b$ admits at most one solution in each of the two sets $S_{00}\cup S_{11}$ and $S_{01}\cup S_{10}$. Consequently, the condition $D_{00}(b)\leq 1$ follows directly from their argument. ∎
3.2 Differential Spectrum when $r=3$
In this subsection, we study the differential spectrum of $F_{3}$ when $q\equiv 11\pmod{12}$. We wish to compute the number of solutions of
$b=F_{3}(x+1)-F_{3}(x)=(x+1)^{3}(1+\chi(x+1))-x^{3}(1+\chi(x)).$ (9)
We already observe that $\delta_{F_{3}}(1,0)=\frac{q+1}{4}$, in Section 3.1. By Lemma 3.2 and 3.4, we have
$\delta_{F_{r}}(1,b)=\begin{cases}D_{00}(b)+1&\text{ if }b=2,\\ D_{00}(b)+D_{01}(b)+D_{10}(b)&\text{ otherwise, }\end{cases}$ (10)
for $b\in\mathbb{F}_{q}^{*}$. We first characterize $D_{00}(b)$, $D_{01}(b)$ and $D_{10}(b)$.
Lemma 3.6.
Let $r=3$ and $q\equiv 3\pmod{4}$. Then,
$D_{00}(b)=\begin{cases}1,&\text{ if }\chi(2b-1)=\chi(2(b-2))=\chi(3),\\ 0,&\text{ otherwise.}\end{cases}$
Proof.
If $x\in S_{00}$, then (9) becomes $\frac{b}{2}=(x+1)^{3}-x^{3}=3x^{2}+3x+1$, so we get
$x^{2}+x-\frac{b-2}{6}=0.$ (11)
The discriminant of (11) is $\frac{2b-1}{3}$. Hence, (11) has two solutions if and only if $\chi(2b-1)=\chi(3)$. If (11) has two solutions $x_{1}$ and $x_{2}$, then $x_{1}x_{2}=-\frac{b-2}{6}$. Since $x\in S_{00}$, we have $\chi(x(x+1))=\chi\left(\frac{b-2}{6}\right)=1$. Since $\chi(x_{1}x_{2})=\chi\left(-\frac{b-2}{6}\right)=-1$, (11) has at most one solution in $S_{00}$. Without loss of generality, we assume that $\chi(x_{1})=1$ and $\chi(x_{2})=-1$. Since $\chi(6(b-2))=\chi(x_{1}(x_{1}+1))=1$ from (11), we have $\chi(x_{1}+1)=1$ so $x_{1}\in S_{00}$.
If $b=\frac{1}{2}$, then (11) has one solution $x=-\frac{1}{2}$. But, we have $\chi(x+1)=\chi\left(\frac{1}{2}\right)\neq\chi\left(-\frac{1}{2}\right)=\chi(x)$ in this case, and hence $\frac{1}{2}\not\in S_{00}$. So, $D_{00}\left(-\frac{1}{2}\right)=0$. ∎
Next, we characterize all $b$ satisfying $D_{01}(b)+D_{10}(b)=1$. Note that we have already shown in Lemma 3.5 that $D_{01}(b)\leq 1$ and $D_{10}(b)\leq 1$.
Lemma 3.7.
Let $r=3$ and $q\equiv 11\pmod{12}$. Then, $D_{01}(b)+D_{10}(b)=1$ if and only if $\chi\left(b\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1$, where $\alpha=2^{\frac{2q-1}{3}}$.
Proof.
If $x\in S_{01}$, then (9) becomes
$x^{3}=-\frac{b}{2}.$ (12)
Since $x\in S_{01}$,
$1=\chi(x)=\chi(x^{3})=\chi\left(-\frac{b}{2}\right)$ (13)
implies that $\chi(2b)=-1$. Applying $\frac{2q-1}{3}$-th power on the both sides of (12), we have
$x=\left(-\frac{b}{2}\right)^{\frac{2q-1}{3}}=-\left(\frac{b}{2}\right)^{\frac{2q-1}{3}}.$
Since $\frac{2q-1}{3}$ is odd, we have $\chi(x)=\chi\left(-\frac{b}{2}\right)=-\chi(2b)=1$. Moreover,
$\chi(x+1)=\chi\left(-\left(\frac{b}{2}\right)^{\frac{2q-1}{3}}+1\right)=\chi\left(-\frac{b^{\frac{2q-1}{3}}-2^{\frac{2q-1}{3}}}{2^{\frac{2q-1}{3}}}\right)=\chi\left(-\frac{b^{\frac{2q-1}{3}}-2^{\frac{2q-1}{3}}}{2}\right)=-1$
if and only if $\chi\left(-2\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1$. Therefore, we have $D_{01}(b)=1$ if and only if
$\chi(2b)=\chi\left(-2\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1.$
We can similarly show that $D_{10}(b)=1$ if and only if
$\chi(2b)=1\text{ and }\chi\left(2\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1.$
We can observe that $D_{01}(b)=1$ and $D_{10}(b)=1$ cannot occur simultaneously, and $D_{01}(b)+D_{10}(b)=1$ if and only if
$1=\chi(2b)\chi\left(2\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=\chi\left(b\left(b^{\frac{2q-1}{3}}-\alpha\right)\right),$
which completes the proof. ∎
For any function $F$ satisfying (4), the differential spectrum of $F$ is defined to be the multiset $DS_{F}=\{\omega_{i}:0\leq i\leq\delta_{F}\}$, where
$\omega_{i}=\#\{b\in\mathbb{F}_{q}:\delta_{F}(1,b)=i\}.$
The following identity for the differential spectrum is well-known :
$\sum_{i=0}^{\delta_{F}}\omega_{i}=\sum_{i=0}^{\delta_{F}}i\cdot\omega_{i}=q$ (14)
Next, we prove the main theorem of this subsection.
Theorem 3.8.
Let $q\equiv 11\pmod{12}$ with $q>11$. Then, $F_{3}(x)=x^{3}+x^{\frac{q+5}{2}}$ is locally-APN. Furthermore, the differential spectrum of $F_{3}$ is given by
$DS_{F_{3}}=\left\{\omega_{0}=\frac{3}{8}(q-3)-\frac{\Gamma}{2},\ \omega_{1}=\frac{1}{2}(q+1)+\Gamma,\ \omega_{2}=\frac{1}{8}(q-3)-\frac{\Gamma}{2},\ \omega_{\frac{q+1}{4}}=1\right\},$
where $\Gamma=\displaystyle\sum_{x\in S}\chi(x)\chi(x-\alpha)$ with $\alpha=2^{\frac{2q-1}{3}}$ and $S=\left\{x\in\mathbb{F}_{q}:\chi(2x^{3}-1)=\chi(2x^{3}-4)=1\right\}$.
Proof.
Combining (8) with Lemma 3.3, we can see that $\delta_{F_{3}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$.
Next, we count the number of $b\in\mathbb{F}_{q}^{*}$ satisfying $\delta_{F_{3}}(1,b)=2$. By Lemma 3.6, we have $D_{00}(2)=0$, and hence $\delta_{F_{3}}(1,2)=1$, using (10). Hence, $\delta_{F_{3}}(1,b)=2$ if and only if $D_{00}(b)=1$ and $D_{01}(b)+D_{10}(b)=1$. So, by Lemmas 3.6 and 3.7, $\delta_{F_{3}}(1,b)=2$ if and only if $\chi(2b-1)=\chi(2b-4)=1$ and $\chi\left(b\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1$. Hence, we have
$8\omega_{2}=$ $-2+\displaystyle\sum_{x\in\mathbb{F}_{q}}(1+\chi(2x-1))(1+\chi(2x-4))(1-\chi\left(x\left(x^{\frac{q+1}{6}}-\alpha\right)\right))$
$=$ $-2+D_{1}-D_{2},$
where
$D_{1}$ $=\displaystyle\sum_{x\in\mathbb{F}_{q}}(1+\chi(2x-1))(1+\chi(2x-4)),$
$D_{2}$ $=\displaystyle\sum_{x\in\mathbb{F}_{q}}(1+\chi(2x-1))(1+\chi(2x-4))\chi\left(x\left(x^{\frac{2q-1}{3}}-\alpha\right)\right).$
By Lemma 2.4,
$D_{1}=\displaystyle\sum_{x\in\mathbb{F}_{q}}(1+\chi(2x-1))(1+\chi(2x-4))=q+\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(2x-1)+\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(2x-4)+\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi((2x-1)(2x-4))=q-1.$
Observe that
$D_{2}$ $=\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x)\chi\left(x^{\frac{2q-1}{3}}-\alpha\right)(1+\chi(2x-1))(1+\chi(2x-4))$
$\underset{x=u^{3}}{=}\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u^{3})\chi\left(u-\alpha\right)(1+\chi(2u^{3}-1))(1+\chi(2u^{3}-4))$
$=\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)(1+\chi(2u^{3}-1))(1+\chi(2u^{3}-4))=4\displaystyle\sum_{u\in S}\chi(u)\chi\left(u-\alpha\right).$
where $S=\left\{x\in\mathbb{F}_{q}:\chi(2x^{3}-1)=\chi(2x^{3}-4)=1\right\}$. Thus,
$\omega_{2}=\frac{1}{8}\left(-2+D_{1}-D_{2}\right)=\frac{1}{8}\left(q-3-4\displaystyle\sum_{x\in S}\chi(x)\chi(x-\alpha)\right).$
We get $\omega_{0}=\frac{3}{8}(q-3)-\frac{\Gamma}{2}$ and $\omega_{1}=\frac{1}{2}(q+1)+\Gamma$ from (14) with the above and $\omega_{\frac{q+1}{4}}=1$.
Since
$D_{2}$ $=4\displaystyle\sum_{u\in S}\chi(u)\chi\left(u-\alpha\right)=\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)(1+\chi(2u^{3}-1))(1+\chi(2u^{3}-4))$
$=-1+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)\chi(2u^{3}-1)+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)\chi(2u^{3}-4)$
$+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)\chi(2u^{3}-1)\chi(2u^{3}-4),$
$\omega_{2}\geq\frac{1}{8}\left(q-2-15\sqrt{q}\right)$, by Lemma 2.6. Hence, we can see that $\omega_{2}>0$, and hence $F_{3}$ is locally-APN if $q>231$, via SageMath. We confirm that $F_{3}$ is also locally-APN for $q<231$ via SageMath (see Table 4). ∎
Note that the above theorem is confirmed via SageMath when $11\leq q<100000$. Table 4 describes the differential spectrum of $F_{3}$ and $\Gamma$ in Theorem 3.8, computed via SageMath, when $q<231$.
$q$ $\Gamma$ $DS_{F_{3}}$
$11$ 22 $\{\omega_{0}=2,\ \omega_{1}=8,\ \omega_{3}=1\}$
$23$ 3-3 $\{\omega_{0}=9,\ \omega_{1}=9,\ \omega_{2}=4,\ \omega_{6}=1\}$
$47$ 5-5 $\{\omega_{0}=19,\ \omega_{1}=19,\ \omega_{2}=8,\ \omega_{12}=1\}$
$59$ 22 $\{\omega_{0}=20,\ \omega_{1}=32,\ \omega_{2}=6,\ \omega_{15}=1\}$
$71$ 55 $\{\omega_{0}=23,\ \omega_{1}=41,\ \omega_{2}=6,\ \omega_{18}=1\}$
$83$ 4-4 $\{\omega_{0}=32,\ \omega_{1}=38,\ \omega_{2}=12,\ \omega_{21}=1\}$
$107$ 2-2 $\{\omega_{0}=40,\ \omega_{1}=52,\ \omega_{2}=14,\ \omega_{27}=1\}$
$131$ 4-4 $\{\omega_{0}=50,\ \omega_{1}=62,\ \omega_{2}=18,\ \omega_{33}=1\}$
$167$ $13$ $\{\omega_{0}=55,\ \omega_{1}=97,\ \omega_{2}=14,\ \omega_{42}=1\}$
$179$ 88 $\{\omega_{0}=62,\ \omega_{1}=98,\ \omega_{2}=18,\ \omega_{45}=1\}$
$191$ 33 $\{\omega_{0}=69,\ \omega_{1}=99,\ \omega_{2}=22,\ \omega_{48}=1\}$
$227$ 0 $\{\omega_{0}=84,\ \omega_{1}=114,\ \omega_{2}=28,\ \omega_{57}=1\}$
Table 4: Differential spectrum $DS_{F_{3}}$ when $q<231$ with $q\equiv 11\pmod{12}$.
3.3 Differential Spectrum when $r=\frac{2q-1}{3}$
In this subsection, we study the differential spectrum of $F_{\frac{2q-1}{3}}$ when $q\equiv 11\pmod{12}$. We wish to compute the number of solutions of
$b=F_{\frac{2q-1}{3}}(x+1)-F_{\frac{2q-1}{3}}(x)=(x+1)^{\frac{2q-1}{3}}(1+\chi(x+1))-x^{\frac{2q-1}{3}}(1+\chi(x)).$ (15)
We already observe that $\delta_{F_{\frac{2q-1}{3}}}(1,0)=\frac{q+1}{4}$, in Section 3.1. As in Section 3.2, we derive the differential spectrum using (10). We first characterize $D_{00}(b)$, $D_{01}(b)$ and $D_{10}(b)$.
Lemma 3.9.
Let $r=\frac{2q-1}{3}$ and $q\equiv 3\pmod{4}$. Then,
$D_{00}(b)=\begin{cases}1&\text{ if }\chi(b(32-b^{3}))=1,\ \chi(b(b^{3}-8))=-1,\\ 0&\text{ otherwise.}\end{cases}$
Proof.
If there exists $x\in S_{00}$ satisfying (15), then (15) leads to
$(x+1)^{\frac{2q-1}{3}}=x^{\frac{2q-1}{3}}+\frac{b}{2}.$
Raising both sides of the above equation to the third power yields
$0=12bX^{2}+6b^{2}X+b^{3}-8,$
where $X=x^{\frac{2q-1}{3}}$. The above equation has two solutions if and only if
$(6b^{2})^{2}-4\times 12b(b^{3}-8)=-12b^{4}+384b=12b(32-b^{3})$
is a square. If then, we obtain
$X=\frac{-6b^{2}\pm(12b(32-b^{3}))^{\frac{q+1}{4}}}{24b}=\frac{-3b^{2}\pm\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}}{12b}$
Then, one has
$x=X^{3}=\frac{\left(-3b^{2}\pm\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}\right)^{3}}{1728b^{3}}=\frac{-36b^{2}\pm(4+b^{3})\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}}{72b^{2}}.$
Hence, we can see that if we denote $x=x_{1}$ is a solution of (15), then the other is $x=x_{2}=-x_{1}-1$. If both $x_{1}$ and $x_{2}$ are in $S_{00}$, then $\chi(x_{1}+1)=\chi(-x_{2})=-1$ and $\chi(x_{2}+1)=\chi(-x_{1})=-1$, a contradiction. Hence, (15) has at most one solution in $S_{00}$. Also, one can check that
$x(x+1)=-\frac{(b^{3}-8)^{3}}{2^{6}\cdot 3^{3}\cdot b^{3}}$
and $\chi(3)=1$, and therefore we have $\chi(x(x+1))=1$ if and only if $\chi(b(b^{3}-8))=-1$. ∎
Next, we characterize all $b$ satisfying $D_{01}(b)+D_{10}(b)=1$. Note that we have already shown in Lemma 3.5 that $D_{01}(b)\leq 1$ and $D_{10}(b)\leq 1$.
Lemma 3.10.
Let $r=\frac{2q-1}{3}$ and $q\equiv 11\pmod{12}$. Then,
$D_{01}(b)+D_{10}(b)=\begin{cases}1&\text{ if }\chi(b(b^{3}-8))=-1,\\ 0&\text{ otherwise.}\end{cases}$
Proof.
If $x\in S_{01}$, then (15) reduces to
$x^{\frac{2q-1}{3}}=-\frac{b}{2}.$ (16)
From the above equation, we get
$\chi\left(\frac{b}{2}\right)=\chi\left(-x^{\frac{2q-1}{3}}\right)=-\chi(x)^{\frac{2q-1}{3}}=-1.$
Raising both sides of (16) to the third power yields
$x=-\frac{b^{3}}{8}.$
Since $x\in S_{01}$, we have
$1=\chi(-(x+1))=\chi\left(\frac{b^{3}-8}{8}\right)=\chi(2(b^{3}-8)).$
Hence, $D_{01}(b)=1$ if and only if $\chi(2b)=-1$ and $\chi(2(b^{3}-8))=1$. Similarly, $D_{10}(b)=1$ if and only if $\chi(2b)=1$ and $\chi(2(b^{3}-8))=-1$. Thus, $\chi(2b)\chi(2(b^{3}-8))=-1$, or equivalently $\chi(b(b^{3}-8))=-1$, which yields the desired result. ∎
Next, we prove the main theorem of this subsection.
Theorem 3.11.
Let $q\equiv 11\pmod{12}$ with $q>11$. Then, $F_{\frac{2q-1}{3}}(x)=x^{\frac{2q-1}{3}}+x^{\frac{q+1}{6}}$ is locally-APN. Furthermore, the differential spectrum of $F_{\frac{2q-1}{3}}$ is given by
$DS_{F_{\frac{2q-1}{3}}}=\left\{\omega_{0}=\frac{q-3}{2},\ \omega_{1}=\frac{q+5}{4},\ \omega_{2}=\frac{q-3}{4},\ \omega_{\frac{q+1}{4}}=1\right\}.$
Proof.
Combining (10) with Lemma 3.9 and Lemma 3.10, we obtain that $\delta(1,b)=2$ if and only if $\chi(b(32-b^{3}))=1$ and $\chi(b(b^{3}-8))=-1$. Hence, we have
$\omega_{2}$ $=\frac{1}{4}\left(\displaystyle\sum_{x\in\mathbb{F}_{q}}\left(1+\chi(x(32-x^{3}))\right)\left(1-\chi(x(x^{3}-8))\right)-3\right)$
$=\frac{1}{4}\left(\displaystyle\sum_{x\in\mathbb{F}_{q}}\left(1+\chi(x(32-x))\right)\left(1-\chi(x(x-8))\right)-3\right)$
$=\frac{1}{4}\left(q-3+\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x(32-x))-\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x(x-8))-\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x^{2}(x-8)(32-x))\right)$
$=\frac{q-3}{4}$
Applying $\omega_{\frac{q+1}{4}}=1$ and $\omega_{2}=\frac{q-3}{4}$ on (14), we have $\omega_{1}=\frac{q+5}{4}$ and $\omega_{0}=\frac{q-3}{2}$. We complete the proof. ∎
We note that the differential spectrum in this case coincides with that of $F_{\frac{q+1}{4}}$ when $q\equiv 7\pmod{8}$ in [10], although the two cases arise under different arithmetic conditions on $q$. This theorem has also been verified numerically using SageMath for all $q$ with $11\leq q<100000$.
4 Boomerang Properties
In this subsection, we study boomerang properties of $F_{r}$. We consider to find the number of common solutions $(x,y)$ of the following system.
$\begin{cases}x^{r}(1+\chi(x))-y^{r}(1+\chi(y))=b,\\ (x+1)^{r}(1+\chi(x+1))-(y+1)^{r}(1+\chi(y+1))=b.\end{cases}$ (17)
Denote by $B_{ijkl}(b)$ the number of solutions of (17) in $S_{ij}\times S_{kl}$, where $i,j,k,l\in\{0,1\}$.
4.1 Boomerang Uniformity of $F_{r}$
In this subsection, we prove the following.
Theorem 4.1.
If $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $\beta_{F_{r}}\leq 2$. In particular, for each parameter set $(p,n,r)$ satisfying the corresponding conditions in Table 3, $\beta_{F_{r}}\leq 2$.
For the exponents $r$ listed in Table 3, we have already shown in Section 3 that $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$. Therefore, it suffices to prove the first assertion of the above theorem. To prove the above theorem, we investigate each $B_{ijkl}(b)$.
Lemma 4.2.
Let $b\in\mathbb{F}_{q}^{*}$. If $\gcd(r,q-1)\in\{1,2\}$, then there is no solution $(x,y)$ of (17) satisfying $x\in(S_{11}\cup\{-1\})$ or $y\in(S_{11}\cup\{-1\})$.
Proof.
By symmetry, we only show that there is no solution $(x,y)$ satisfying $x\in(S_{11}\cup\{-1\})$. Suppose that there is such a solution $(x,y)$ satisfying $x\in(S_{11}\cup\{-1\})$. If $y\in S_{00}$, then (17) reduces to $y^{r}=(y+1)^{r}=-\frac{b}{2}$. This impies that $y^{\gcd(r,q-1)}=(y+1)^{\gcd(r,q-1)}$. If $\gcd(r,q-1)=1$, then we have $y=y+1$, a contradiction. If $\gcd(r,q-1)=2$, then we obtain $y^{2}=(y+1)^{2}$, which implies $y=-\frac{1}{2}$. Then, we get $\chi(y)=\chi\left(-\frac{1}{2}\right)=\chi\left(\frac{1}{2}\right)=\chi(y+1)=1$, a contradiction.
If $y\not\in S_{00}$, then at least one of equations in (17) leads to $b=0$, a contradiction. ∎
Lemma 4.3.
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}^{*}$. Then, $B_{0101}(b)=B_{0110}(b)=B_{1010}(b)=B_{1001}(b)=0$.
Proof.
The proof is very similar with Lemma 3.8 of our previous paper [10], and we omit here. ∎
Lemma 4.4.
Let $b\in\mathbb{F}_{q}^{*}$. If
$\begin{cases}x^{r}-y^{r}=\frac{b}{2},\\ (x+1)^{r}-(y+1)^{r}=\frac{b}{2}.\end{cases}$ (18)
does not have a solution $(x,y)\in S_{00}\times S_{00}$, then $B_{0000}(b)=0$. Furthermore, if $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$, then $B_{0000}(b)=0$.
Proof.
If $(x,y)\in S_{00}\times S_{00}$, then (17) leads to (18). Then we have $x^{r}-y^{r}=\frac{b}{2}=(x+1)^{r}-(y+1)^{r}$, which implies $(x+1)^{r}-x^{r}=(y+1)^{r}-y^{r}$. By Lemma 3.3, $D_{00}(0)=0$ and hence $(x+1)^{r}-x^{r}=(y+1)^{r}-y^{r}=c$ for some $c\in\mathbb{F}_{q}^{*}$. Since $D_{00}(c)\leq 1$ for all $c\in\mathbb{F}_{q}^{*}$, we have $x=y$, and hence $b=0$, a contradiction. ∎
By applying Lemma 4.4 to the results of Theorem 3.1 and Lemma 3.6, we obtain that $B_{0000}(b)=0$ for every nonzero $b\in\mathbb{F}_{q}$ for all $r$ listed in Table 3.
Lemma 4.5.
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}^{*}$. Then, $B_{0100}(b)=B_{1000}(b)=0$ or $B_{0001}(b)=B_{0010}(b)=0$.
Proof.
Assume that $B_{0100}(b)>0$. Then, (17) reduces to
$\begin{cases}x^{r}-y^{r}=\frac{b}{2},\\ -(y+1)^{r}=\frac{b}{2}.\end{cases}$
From the second equation, we have $\chi\left(\frac{b}{2}\right)=-1$.
Suppose that $B_{0001}(b)>0$. Then, the second equation of (17) reduces to $(x+1)^{r}=\frac{b}{2}$, and hence there is $x\in S_{00}$ such that $(x+1)^{r}=\frac{b}{2}$. Then we have $\chi\left((x+1)^{r}\right)=1=-1=\chi\left(\frac{b}{2}\right)$, a contradiction. Thus, $B_{0001}(b)=0$. Similarly, if $B_{0010}(b)>0$, then the first equation of (17) reduces to $x^{r}=\frac{b}{2}$. Then, we obtain $\chi\left(x^{r}\right)=1=-1=\chi\left(\frac{b}{2}\right)$, a contradiction, and hence $B_{0010}(b)=0$.
Similarly, if $B_{1000}(b)>0$, then we have $\chi\left(\frac{b}{2}\right)=\chi(-y^{r})=-1$, and hence $B_{0001}(b)=B_{0010}(b)=0$.
By symmetry, if $B_{0001}(b)>0$ or $B_{1000}(b)>0$, then $B_{0100}(b)=B_{1000}(b)=0$. ∎
Lemma 4.6.
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}^{*}$. If $\gcd(r,q-1)\in\{1,2\}$, then $B_{0001}(b)\leq 1$, $B_{0010}(b)\leq 1$, $B_{0100}(b)\leq 1$, and $B_{1000}(b)\leq 1$.
Proof.
We only give the proof of $B_{0001}(b)\leq 1$, because proofs of the other cases are similar. If $(x,y)\in S_{00}\times S_{01}$, then equations in (17) are reduced to
$\begin{cases}x^{r}-y^{r}=\frac{b}{2},\\ (x+1)^{r}=\frac{b}{2}.\end{cases}$ (19)
The second equation of (19) has $\gcd(r,q-1)$ solutions. If $\gcd(r,q-1)=1$, then there exists a unique $x\in\mathbb{F}_{q}$ satisfying it. If $\gcd(r,q-1)=2$, then $r$ is even and hence $x^{r}=(-x)^{r}$ for all $x\in\mathbb{F}_{q}$. Thus, if $x=\theta$ is a solution of $(x+1)^{p^{k}+1}=\frac{b}{2}$, the other solution is $x=-\theta-2$. If both $\theta$ and $\theta-2$ lie in $S_{00}$, then $1=\chi(\theta+1)=\chi(-\theta-1)=-\chi(\theta+1)$, a contradiction. Therefore, the second equation of (19) admits at most one solution in $S_{00}$ in either case. For such $x$, the first equation of (19) has at most one solution $y$ with $\chi(y)=1$ in the same manner. Hence, $B_{0001}(b)\leq 1$, which completes the proof. ∎
Now, we are ready to prove the main theorem of this subsection.
Proof of Theorem 4.1.
It remains to consider the case where (17) has a solution with $x=0$ or $y=0$, when $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$. We only consider the case where (17) has a solution with $x=0$, since the case $y=0$ can be proved similarly by symmetry. In this case, (17) leads to
$(1+\chi(y))y^{r}=2-(1+\chi(y+1))(y+1)^{r}=b.$ (20)
If (20) has a solution $y\in\mathbb{F}_{q}$ with $\chi(y)=-1$ or $y=0$, then we have $b=0$, a contradiction. If (20) has a solution $y\in S_{00}$, then we have $2y^{r}=2-2(y+1)^{r}=b$, so $y\in S_{00}$ is a solution of $(y+1)^{r}-y^{r}=1$. Since we already see that $D_{00}(1)\leq 1$, there exists at most one $y\in S_{00}$ satisfying $2y^{r}=2-2(y+1)^{r}=b$. Let $y=\theta\in S_{00}$ be the solution if it exists. Note that $\chi\left(\frac{b}{2}\right)=\chi(-\theta^{r})=-1$. By Lemmas 4.2, 4.3 and 4.4,
$\beta_{F_{r}}(1,b)\leq 1+B_{0001}(b)+B_{0010}(b)+B_{0100}(b)+B_{1000}(b).$ (21)
If there is another solution of (17) with $x\in S_{00}$, then we have $x^{r}=\frac{b}{2}$ or $(x+1)^{r}=\frac{b}{2}$. If $\gcd(r,q-1)=1$, then we have $x=-\theta$ or $x=-\theta-1$, respectively. But, we observe that $\theta,-\theta-1\not\in S_{00}$, because $\chi(-\theta)=-1$. If $\gcd(r,q-1)=2$, there is no such $x\in\mathbb{F}_{q}$ because $r$ is even and $\chi\left(\frac{b}{2}\right)=-1$. Hence, $B_{0001}(b)=B_{0010}(b)=0$. If there is another solution of (17) in $S_{10}\times S_{00}$, then $y^{r}=-\frac{b}{2}$. If $\gcd(r,q-1)=1$, then $y=\theta$. Then, $(x+1)^{r}=(\theta+1)^{r}+\frac{b}{2}=1$, and hence $x=0$, a contradiction. If $\gcd(r,q-1)=2$, then two solutions of $y^{r}=-\frac{b}{2}$ is $y=\pm\theta$, and hence $y=\theta$ since $\chi(-\theta)=-1$. Then, we have $(x+1)^{r}=1$ which implies that $(x+1)^{2}=1$. Since $x\neq 0$, we obtain x=2x=-2. Then, $\chi(x+1)=\chi(-1)=-1$, which is a contradiction to $x\in S_{10}$, and hence $B_{1000}(b)=0$. By Lemma 4.6 with (21), we obtain $\beta_{F_{r}}(1,b)\leq 2$.
If (20) has a solution $y\in S_{01}$, then we have $2y^{r}=2=b$, which applies $y^{r}=-1$ and hence $r$ is odd. Then we obtain $\gcd(r,q-1)=1$ and y=1y=-1, which contradicts that $y\in S_{01}$.
If there is no solution of (17) with $x=0$ or $y=0$, then by Lemmas 4.2, 4.3 and 4.4,
$\beta_{F_{r}}(1,b)=B_{0001}(b)+B_{0010}(b)+B_{0100}(b)+B_{1000}(b).$
Combining Lemma 4.5 and Lemma 4.6 with the above equation, we have $\beta_{F}(1,b)\leq 2$, which completes the proof. ∎
4.2 Boomerang Spectrum of $F_{2}$ when $p=3$
In the above subsection, we showed that the boomerang uniformity of $F_{r}$ is at most 2 for several values of $r$, especially when $r=\frac{3^{k}+1}{2}$, that is, generalized Coulter-Matthews exponent. Next, we consider the special case $k=1$ and prove that the boomerang uniformity of $F_{r}$ is actually $1$ in this case. When $k=1$, we have $r=2$, and the boomerang uniformity of $F_{2}$ was studied in [14]. Theorem $8$ of [14] proved that the boomerang uniformity of $F_{2}$ equals $2$ for $q\geq 9613^{2}$, and conjectured that this holds for $q>307$. However, we demonstrate that this result does not hold when $p=3$ ; specifically, the boomerang uniformity is $1$ in this case. Moreover, we also provide the full boomerang spectrum of $F_{2}$.
By Lemma 18 of [14], (17) with $r=2$ has no solution with $x=0$ or $y=0$. Hence, as observed in the above subsection, $\beta_{F_{2}}(1,b)=B_{0001}(b)+B_{0010}(b)+B_{0100}(b)+B_{1000}(b)$ holds, and that $B_{0001}(b),B_{0010}(b),B_{0100}(b),B_{1000}(b)\leq 1$. We next characterize the conditions under which each of these quantities equals 1 and show that no two of these conditions can hold simultaneously.
Lemma 4.7.
Let $p=3$, $n$ be odd, $r=2$, and $b\in\mathbb{F}_{q}^{*}$. Then,
  • $B_{0001}(b)=1$ if and only if
    $\chi(b)=\chi\left(b^{\frac{q+1}{4}}+1\right)=\chi\left(b^{\frac{q+1}{4}}-1\right)=\chi\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}+1\right)=-1.$ (22)
  • $B_{0010}(b)=1$ if and only if
    $\chi(b)=\chi\!\left(b^{\frac{q+1}{4}}-1\right)=\chi\!\left(\left(b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)=-1,\ \chi\!\left(b^{\frac{q+1}{4}}+1\right)=1.$ (23)
  • $B_{0100}(b)=1$ if and only if
    $\chi(b)=\chi\!\left(b^{\frac{q+1}{4}}-1\right)=\chi\!\left(b^{\frac{q+1}{4}}+1\right)=1,\ \chi\!\left(\left(b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}+1\right)=-1.$ (24)
  • $B_{1000}(b)=1$ if and only if
    $\chi(b)=\chi\!\left(b^{\frac{q+1}{4}}+1\right)=1,\ \chi\!\left(b^{\frac{q+1}{4}}-1\right)=\chi\!\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)=-1.$ (25)
Proof.
If there is a solution $(x,y)\in S_{00}\times S_{01}$ in (17) with $r=2$, then (17) reduces to
$\begin{cases}x^{2}-y^{2}=-b,\\ (x+1)^{2}=-b.\end{cases}$ (26)
Raising both sides of the second equation in (26) to the $\frac{q+1}{4}$-th power yields $x+1=-b^{\frac{q+1}{4}}$ and $\chi(b)=-1$, since $\frac{q+1}{4}$ is odd when $q=3^{n}$ with $n$ odd and $\chi(x+1)=1$. Substituting $x=-b^{\frac{q+1}{4}}-1$ into the first equation in (26), we obtain
$y^{2}=x^{2}+b=\left(-b^{\frac{q+1}{4}}-1\right)^{2}+b=-b^{\frac{q+1}{4}}+1,$
and hence $y=\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}$. Consequently, existence of solution $(x,y)\in S_{00}\times S_{01}$ satisfying (26) implies that $x$ and $y$ are uniquely determined. Furthermore, it is straightforward to check
$\chi(b)=\chi\!\left(b^{\frac{q+1}{4}}+1\right)=\chi\!\left(b^{\frac{q+1}{4}}-1\right)=\chi\!\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}+1\right)=-1.$
By symmetry, proofs for the other cases are similar and we omit them.
Conversely, if conditions in (22) are satisfied, then we can easily see that
$(x,y)=\left(-b^{\frac{q+1}{4}}-1,\ \left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}\right)$
is a solution of (26), by the above process. ∎
Applying Lemma 2.2, the boomerang spectrum of $F$ is defined to be the multiset $BS_{F}=\{\nu_{i}:0\leq i\leq\beta_{F}\}$, where
$\nu_{i}=\#\{b\in\mathbb{F}_{p^{n}}^{*}:\beta_{F}(1,b)=i\}.$
The following identity for the boomerang spectrum is well-known :
$\sum_{i=0}^{\beta_{F}}\nu_{i}=q-1.$ (27)
Theorem 4.8.
Let $p=3$ and $n\geq 3$ odd. Then $\beta_{F_{2}}=1$. Furthermore, the boomerang spectrum of $F_{2}$ is given by
$BS_{F_{2}}=\left\{\nu_{0}=\frac{3q-5+2\Lambda}{4},\quad\nu_{1}=\frac{q+1-2\Lambda}{4}\right\},$
where $\Lambda=\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x+1)\chi(x^{2}+1)$.
Proof.
Since any two of (22), (23), (24), and (25) in Lemma 4.7 cannot occur simultaneously, we obtain $\beta_{F_{2}}\leq 1$.
If $\chi\left(b^{\frac{q+1}{4}}-1\right)=-1$, then
$\chi\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}+1\right)\chi\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)=\chi\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{2}}-1\right)$
$=\chi\left(\chi\left(-b^{\frac{q+1}{4}}+1\right)\cdot\left(-b^{\frac{q+1}{4}}+1\right)-1\right)=\chi\left(-b^{\frac{q+1}{4}}\right)=-\chi(b).$
Hence, $B_{0001}(b)=1$ if and only if
$\chi(b)=-1,\ \chi\!\left(b^{\frac{q+1}{4}}+1\right)=-1,\ \chi\!\left(b^{\frac{q+1}{4}}-1\right)=-1,$
$-1=\chi\!\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}+1\right)=-\chi\left(b\right)\chi\!\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)=\chi\!\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right).$
Thus, $B_{0001}(b)+B_{1000}(b)=1$ if and only if
$\chi(b)\chi\left(b^{\frac{q+1}{4}}+1\right)=1,\ \chi\left(b^{\frac{q+1}{4}}-1\right)=\chi\left(\left(-b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)=-1.$ (28)
Similarly, $B_{0100}(b)+B_{0010}(b)=1$ if and only if
$\chi(b)\chi\left(b^{\frac{q+1}{4}}-1\right)=\chi\left(b^{\frac{q+1}{4}}+1\right)=1,\ \chi\left(\left(b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)=-1.$ (29)
(29) coincides with (28) after substituting $b$ by b-b. Hence, $\nu_{1}$ equals twice the number of $b$ satisfying (29), and is therefore given by
$\nu_{1}=\displaystyle\frac{1}{8}(2A-4)=\displaystyle\frac{1}{4}(A-2),$ (30)
where
$A=\displaystyle\sum_{b\in\mathbb{F}_{q}}\left(1+\chi(b)\chi\!\left(b^{\frac{q+1}{4}}-1\right)\right)\left(1+\chi\!\left(b^{\frac{q+1}{4}}+1\right)\right)\left(1-\chi\left(\left(b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)\right).$
Let $y=b^{\frac{q+1}{4}}$. Then, applying Lemma 2.4 and Lemma 2.5,
$A$ $=q+\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi(y(y-1))+\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi(y+1)-\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi\left((y+1)^{\frac{q+1}{4}}-1\right)+\displaystyle\sum_{y\in\mathbb{F}_{q}}\chi(y(y^{2}-1))$
$-\displaystyle\sum_{y\in\mathbb{F}_{q}}\left(\chi(y(y-1))+\chi(y+1)+\chi(y(y^{2}-1))\right)\chi\left((y+1)^{\frac{q+1}{4}}-1\right)$
$=q-1-\displaystyle\sum_{y\in\mathbb{F}_{q}}\left(\chi(y(y-1))+\chi(y+1)+\chi(y(y^{2}-1))\right)\chi\left((y+1)^{\frac{q+1}{4}}-1\right).$
If $t=(y+1)^{\frac{q+1}{4}}$, then
$y=\begin{cases}t^{2}-1&\text{ if }\chi(t)=1,\\ -t^{2}-1&\text{ if }\chi(t)=-1.\end{cases}$
$A$ $=q-2-\displaystyle\sum_{\chi(t)=1}\left(\chi((t^{2}-1)(t^{2}-2))+\chi(t^{2})+\chi(t^{2}(t^{2}-1)(t^{2}-2))\right)\chi\left(t-1\right)$
$-\displaystyle\sum_{\chi(t)=-1}\left(\chi((-t^{2}-1)(-t^{2}-2))+\chi(-t^{2})+\chi(-t^{2}(-t^{2}-1)(-t^{2}-2))\right)\chi\left(t-1\right)$
$=q-2-\displaystyle\sum_{\chi(t)=1}\left(2\chi((t^{2}-1)(t^{2}-2))+1\right)\chi\left(t-1\right)+\displaystyle\sum_{\chi(t)=-1}\chi(t-1).$
Since
$\sum_{\chi(t)=1}\chi(t-1)-\displaystyle\sum_{\chi(t)=-1}\chi(t-1)=\frac{1}{2}\left(\displaystyle\sum_{t\in\mathbb{F}_{q}^{*}}(1+\chi(t))\chi(t-1)-\displaystyle\sum_{t\in\mathbb{F}_{q}^{*}}(1-\chi(t))\chi(t-1)\right)=\displaystyle\sum_{t\in\mathbb{F}_{q}}\chi(t(t-1))=-1,$
we obtain
$A$ $=q-1-2\displaystyle\sum_{\chi(t)=1}\chi((t^{2}-1)(t^{2}-2)(t-1))=q+1-2\displaystyle\sum_{\chi(t)=1}\chi((t+1)(t^{2}-2))$
$=q+1-\displaystyle\sum_{t\in\mathbb{F}_{q}^{*}}(1+\chi(t))\chi((t+1)(t^{2}+1))$
$=q+1-\displaystyle\sum_{t\in\mathbb{F}_{q}^{*}}\chi((t+1)(t^{2}+1))-\displaystyle\sum_{t\in\mathbb{F}_{q}}\chi(t(t+1)(t^{2}+1)).$
Let $z=\frac{t(t+1)}{t^{2}+1}$. Then, $z=1$ if and only if $t=1$. If $t=1$, then $(z-1)t^{2}-t+z=0$. The discriminant of this quadratic equation on $t$ can be computed as $(-1)^{2}-z(z-1)=1+z-z^{2}$. Hence, applying z=s1z=-s-1, we obtain
$\displaystyle\sum_{t\in\mathbb{F}_{q}}\chi(t(t+1)(t^{2}+1))$ $=1+\displaystyle\sum_{t\in\mathbb{F}_{q}\setminus\{1\}}\chi(t(t+1)(t^{2}+1))=1+\displaystyle\sum_{t\in\mathbb{F}_{q}\setminus\{1\}}\chi\left(\displaystyle\frac{t(t+1)}{t^{2}+1}\right)$
$=1+\displaystyle\sum_{z\in\mathbb{F}_{q}\setminus\{1\}}(1+\chi(1+z-z^{2}))\chi(z)=-1+\displaystyle\sum_{z\in\mathbb{F}_{q}}\chi(z)\chi(1+z-z^{2})$
$=-1+\displaystyle\sum_{s\in\mathbb{F}_{q}}\chi(-s-1)\chi(-s^{2}-1)=-1+\displaystyle\sum_{s\in\mathbb{F}_{q}}\chi(s+1)\chi(s^{2}+1)$
Thus, we get
$A$ $=q+2-\displaystyle\sum_{t\in\mathbb{F}_{q}}\chi((t+1)(t^{2}+1))-\left(-1+\displaystyle\sum_{t\in\mathbb{F}_{q}}\chi((t+1)(t^{2}+1))\right)$
$=q+3-2\displaystyle\sum_{t\in\mathbb{F}_{q}}\chi((t+1)(t^{2}+1)).$
Hence, (30) reduces to
$\nu_{1}=\displaystyle\frac{1}{4}\left(A-2\right)=\displaystyle\frac{1}{4}\left(q+1-2\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi((x+1)(x^{2}+1))\right)=\displaystyle\frac{1}{4}(q+1-2\Lambda)$
as desired. Applying (27), we have
$\nu_{0}=q-1-\nu_{1}=\displaystyle\frac{1}{4}\left(3q-5+2\Lambda\right).$
By Lemma 2.6, $|\Lambda|\leq 2\sqrt{q}$. Hence, we have $\nu_{1}=\frac{1}{4}(q+1-2\Lambda)\geq\frac{1}{4}(q+1-4\sqrt{q})$. We confirm that $\frac{1}{4}(q+1-4\sqrt{q})>1$ when $q>22$, and therefore $\beta_{F_{2}}=1$ when $n\geq 3$ is odd. ∎
The boomerang spectrum of $F_{2}$ for $3\leq n\leq 9$, verified using SageMath, is shown in Table 5.
$n$ $\Lambda$ $BS_{F_{2}}$
3 10-10 $\{\nu_{0}=14,\ \nu_{1}=12\}$
5 22 $\{\nu_{0}=182,\ \nu_{1}=60\}$
7 $86$ $\{\nu_{0}=1682,\ \nu_{1}=504\}$
9 190-190 $\{\nu_{0}=14666,\ \nu_{1}=5016\}$
Table 5: Boomerang spectrum $BS_{F_{2}}$ when $p=3$, $n$ is odd, $3\leq n\leq 9$.
By Lemma 2.8 and Theorem 4.8, the following result follows immediately.
Corollary 4.9.
Let $p=3$ and $n\geq 3$ be odd. Then, $\beta_{F_{\frac{3^{n-1}+1}{2}}}=1$. Furthermore, $F_{\frac{3^{n-1}+1}{2}}$ has the same boomerang spectrum as $F_{2}$, which is given in Theorem 4.8.
5 Conclusion
In this paper, we show that if $\gcd(q-1,r)\mid 2$ and $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$, then $F_{r}$ is locally-APN with boomerang uniformity at most $2$. Moreover, it holds that $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$ for the exponents $r$ listed in Table 3. We further investigate the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$. In addition, by analyzing the boomerang spectrum of $F_{2}$ when $p=3$, we prove that its boomerang uniformity is equal to $1$, which contradicts the result of [14], where it was claimed that the boomerang uniformity of $F_{2}$ equals $2$ for sufficiently large $q$.
It should be noted that Table 3 does not exhaust all APN power functions with $\gcd(r,q-1)=2$. Preliminary experiments suggest that $D_{00}(b)\leq 1$ may also hold for some additional exponents, which we plan to investigate in future work. We also plan to further explore how this idea can be extended in future work.
Acknowledgements : This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2021R1C1C2003888). Soonhak Kwon was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education (No. RS-2019-NR040081).